### v1

@@ Abstract
Dengue, Spatial Epidemiology, Communicability Curvature, Network Science, Bayesian Hierarchical Models, INLA, Urban Connectivity, Spatial Dependence
@@ 1 Introduction
Dengue remains one of the most significant vector-borne diseases worldwide, with estimates indicating nearly 390 million infections annually and recurrent outbreaks across tropical and subtropical regions. Its primary vector, the Aedes aegypti mosquito, exhibits complex ecological and epidemiological dynamics that continue to challenge surveillance and control programs. Despite the extensive body of knowledge on vector biology, uncertainties persist regarding how human cases cluster in time and space, and how these clusters can be translated into transmission trajectories that are informative for public health decision-making.
Traditional epidemiological models, such as compartmental frameworks (SIR and related variants), capture aggregated trends but often overlook the fine-scale spatiotemporal dynamics underlying vector–host interactions. Graph- and topology-based approaches offer a complementary perspective by explicitly representing the potential connections among cases, enabling the investigation of transmission chains, local clusters, and structural properties of outbreaks. Although such approaches have gained traction within network epidemiology, they remain underexplored in the context of dengue transmission in urban environments.
The construction of realistic topological models for dengue depends on defining biologically plausible temporal and spatial windows. Empirical studies provide key parameters: the average flight range of Aedes aegypti is approximately 105 m, rarely exceeding 200 m [1], while the extrinsic incubation period of the dengue virus within the mosquito varies between 6 and 14 days [2]. Incorporating these parameters enables the development of models that approximate either (i) transmission chains, in which one case leads to another through the vector–host cycle, or (ii) co-infection clusters, in which multiple cases share a common infection source within a defined spatiotemporal window.
In this study, we introduce a modeling framework grounded in street-level time series, taking advantage of the fact that public health records in Recife include the precise location of each reported case. This enables us to relate and compare temporal patterns across streets and to identify statistical zones of influence emerging from their time series. We employ a contemporary network metric—communicability curvature—which quantifies the importance of a given edge to the integrity of the graph’s connected component. The rationale behind this analysis is to identify influential “arteries,” or critical corridors, that play a disproportionately important role in the topological structure of dengue transmission. The removal or disruption of such edges would likely hinder the maintenance of dengue incidence, providing a highly localized and cost-effective strategy for outbreak mitigation.
@@ 2.1 Local Time Series Model
This model is closely aligned with recent formulations used in spatial epidemiology and the study of complex systems. Each vertex corresponds to a georeferenced point representing the street where cases were recorded, and each vertex is associated with a local incidence time series. Edges are established between vertices whose temporal similarity—here measured by the correlation between their respective series—exceeds a predefined threshold, indicating epidemiological synchrony between distinct locations.
@@
Incorporating this statistical layer allows the identification of spatiotemporal clusters with greater robustness than purely spatial approaches. Analogous correlation-based graph models were widely explored during the COVID-19 pandemic, where they were used to quantify systemic fragility and epidemic risk [3].
@@
To ensure biological plausibility, an additional spatial pruning criterion is applied: only pairs of points separated by less than 600 m are connected. This threshold represents approximately three times the average flight range of Aedes aegypti [1], accounting for geocoding uncertainty and for the approximation inherent in using street centroids.
@@ 3 Theoretical and Empirical Results
Figure 1: Structural representation of streets and connectivity between neighborhoods via curvature in 2015.
(a) 2015
(b) 2019
(c) 2024
Figure 2: Spatial evolution of mean curvature across neighborhoods of Recife in three critical years (2015, 2019, and 2024).
@@ 3.1 Theoretical Results
The construction of the epidemiological graph allows the communicability $C_{ij}(\beta)$ to be interpreted as a measure of accessibility between vertices, incorporating all possible transmission routes weighted by the topological cost of each path [6, 7]. Before presenting the key results, we introduce the formal definition of communicability and the role of the parameter $\beta$, both essential for the subsequent epidemiological interpretation.
@@
Since $(A^{k})_{ij}$ counts the number of walks of length $k$ between $i$ and $j$, communicability sums all possible routes, weighting longer paths by decaying factors $\beta^{k}/k!$. Thus, $C_{ij}(\beta)$ captures structural accessibility across multiple scales, from local neighborhoods to global connections.
@@ 4.3 Classical Spatial Models (SAR and SAC)
Traditional spatial models—the Spatial Autoregressive Model (SAR) and the Spatial Autoregressive Combined (SAC)—were fitted as an intermediate step to assess residual spatial dependence prior to hierarchical Bayesian modeling. Both models used the Queen contiguity matrix, despite the structural fragmentation of the mesh (33 disconnected components), a typical feature of cities with heterogeneous morphology.
Despite this topological limitation, the models produced stable estimates consistent with the remaining results. The SAR model exhibited Pseudo-$R^{2}=0.424$, while the SAC model yielded a similar value ($0.421$). These values indicate that a substantial portion of spatial variability is captured by the covariates, although enough residual autocorrelation remains to justify a hybrid spatial–random-effects approach (e.g., BYM2 or SPDE–INLA).
@@
The spatial lag term $Wy$, however, was not statistically significant (SAR: $p=0.20$; SAC: $p=0.40$), suggesting that spatial dependence does not operate primarily through autocorrelation between adjacent areas, but rather through structural gradients in the urban environment—supporting the use of hierarchical models such as BYM2, which represent spatial smoothing more flexibly.
@@ 4.4 INLA/BYM2 Model: Latent Spatial Structure and Robustness
As the final stage of the empirical analysis, we estimated a hierarchical spatial model of the Besag–York–Mollié family in its BYM2 parametrization, using the Integrated Nested Laplace Approximation (INLA). This model is widely employed in contemporary spatial epidemiology because it decomposes variability into a structured component (neighborhood-based) and an unstructured component, while providing accurate inference for the latent process via hyperparameters.
@@
Table 2: Fixed effects of the INLA/BYM2 model.
Variable mean sd 2.5% 50% 97.5%
Intercept 23662.47 1819.81 20093.61 23662.43 27231.56
year_num -11.75 0.90 -13.52 -11.75 -9.98
curv -46.41 3.99 -54.24 -46.41 -38.58
year_mean_precip -0.02 0.01 -0.04 -0.02 -0.004
precipitation_lag3_mm 0.37 0.07 0.24 0.37 0.50
rainy_season 21.58 3.32 15.08 21.58 28.09
slum_proportion -25.84 12.93 -51.19 -25.85 -0.48
population_density 0.0056 0.0004 0.0048 0.0056 0.0065
richest_income_share -28.43 8.34 -44.79 -28.43 -12.06
total_population -0.00044 0.00015 -0.00074 -0.00044 -0.00014
@@
Table 3: BYM2 hyperparameters.
Hyperparameter mean sd 2.5% 50% 97.5%
Precision (Gaussian obs.) 1.39e-04 5.77e-06 1.28e-04 1.39e-04 1.51e-04
@@
Phi (structured proportion) 5.41e-10 1.16e-10 3.02e-10 5.48e-10 7.26e-10
@@
These results indicate that the geometry of the urban network—not mere geographic proximity—structures dengue diffusion in Recife.

### v2

@@ Abstract
Dengue, Spatial Epidemiology, Correlation-Weighted Communicability Curvature, Network Science, Bayesian Hierarchical Models, INLA
[…]
[1] organization=Núcleo de Formação de Docentes, Universidade Federal de Pernambuco (UFPE), addressline=, city=Caruaru, postcode=, state=PE, country=Brazil
[…]
[2] organization=Núcleo Interdisciplinar de Ciências Exatas e da Natureza (NICEN), Universidade Federal de Pernambuco (UFPE), addressline=, city=Caruaru, postcode=, state=PE, country=Brazil
[…]
[3] organization=Departamento de Engenharia Cartográfica, Universidade Federal de Pernambuco (UFPE), addressline=, city=Recife, postcode=, state=PE, country=Brazil
{graphicalabstract}
{highlights}
We introduce correlation-weighted communicability curvature as a structural metric to explain spatiotemporal dengue patterns in dense urban environments.
Curvature substantially reduces the spatially structured component in the BYM2-INLA model, capturing variability usually attributed to geographic adjacency.
Annual shifts in mean curvature reveal changes in the underlying transmission architecture, with epidemic years showing more negative and structurally integrated networks.
NDWI shows a significant effect, though its interpretation requires caution due to noise introduced by persistent cloud cover in tropical regions.
Results highlight the value of integrating urban networks, spectral graph metrics, and hierarchical Bayesian modeling for high-resolution epidemiological surveillance.
@@ 1 Introduction
Dengue remains one of the most significant vector-borne diseases worldwide, with estimates indicating nearly 390 million infections annually and recurrent outbreaks across tropical and subtropical regions. Its primary vector, the Aedes aegypti mosquito, exhibits complex ecological and epidemiological dynamics that continue to challenge surveillance and control programs [1, 2, 3].. Despite the extensive body of knowledge on vector biology, uncertainties persist regarding how human cases cluster in time and space, and how these clusters can be translated into transmission trajectories that are informative for public health decision-making.
Traditional epidemiological models, such as compartmental frameworks (SIR and related variants), capture aggregated trends but often overlook the fine-scale spatiotemporal dynamics underlying vector–host interactions [4, 5, 6]. Graph- and topology-based approaches offer a complementary perspective by explicitly representing the potential connections among cases, enabling the investigation of transmission chains, local clusters, and structural properties of outbreaks. Although such approaches have gained traction within network epidemiology [7, 8], they remain underexplored in the context of dengue transmission in urban environments.
The construction of realistic topological models for dengue depends on defining biologically plausible temporal and spatial windows. Empirical studies provide key parameters: the average flight range of Aedes aegypti is approximately 105 m, rarely exceeding 200 m [11], while the extrinsic incubation period of the dengue virus within the mosquito varies between 6 and 14 days [12]. Incorporating these parameters enables the development of models that approximate either (i) transmission chains, in which one case leads to another through the vector–host cycle, or (ii) co-infection clusters, in which multiple cases share a common infection source within a defined spatiotemporal window [9, 10].
In this study, we introduce a modeling framework grounded in street-level time series, taking advantage of the fact that public health records in Recife include the precise location of each reported case. This enables us to relate and compare temporal patterns across streets and to identify statistical zones of influence emerging from their time series. We employ a contemporary network metric — communicability curvature [7, 15] —which quantifies the importance of a given edge to the integrity of the graph’s connected component. The rationale behind this analysis is to identify influential “arteries,” or critical corridors, that play a disproportionately important role in the topological structure of dengue transmission. The removal or disruption of such edges would likely hinder the maintenance of dengue incidence, providing a highly localized and cost-effective strategy for outbreak mitigation.
@@ 2.1 Local Time Series Model
This model is closely aligned with recent formulations used in spatial epidemiology and the study of complex systems [13, 8]. Each vertex corresponds to a georeferenced point representing the street where cases were recorded, and each vertex is associated with a local incidence time series. Edges are established between vertices whose temporal similarity—here measured by the correlation between their respective series—exceeds a predefined threshold, indicating epidemiological synchrony between distinct locations.
To illustrate the resulting structural connectivity, Figure 1 shows the functional street-level graph superimposed on the geographic map of Recife. This visualization highlights the major channels of structural flow induced by correlated dengue activity and provides a geometric understanding of the network used throughout the study.
Figure 1: Street-level functional connectivity graph for Recife. Nodes represent geocoded dengue case locations, and edges denote correlated local incidence time series. The graph is superimposed on the geographic map to reveal the underlying transmission topology, highlighting structural pathways and spatial constraints that shape diffusive spread.
@@
Incorporating this statistical layer allows the identification of spatiotemporal clusters with greater robustness than purely spatial approaches. Analogous correlation-based graph models were widely explored during the COVID-19 pandemic, where they were used to quantify systemic fragility and epidemic risk [14, 24, 5].
@@
To ensure biological plausibility, an additional spatial pruning criterion is applied: only pairs of points separated by less than 600 m are connected. This threshold represents approximately three times the average flight range of Aedes aegypti [11], accounting for geocoding uncertainty and for the approximation inherent in using street centroids [25, 6].
@@ 3 Theoretical and Empirical Results
Figure 2: Mean communicability curvature by neighborhood in 2015.
Figure 3: Mean communicability curvature by neighborhood in 2019.
Figure 4: Mean communicability curvature by neighborhood in 2024.
@@ 3.1 Theoretical Results
The construction of the epidemiological graph allows the communicability $C_{ij}(\beta)$ to be interpreted as a measure of accessibility between vertices, incorporating all possible transmission routes weighted by the topological cost of each path [15, 7, 16, 17]. Before presenting the key results, we introduce the formal definition of communicability and the role of the parameter $\beta$, both essential for the subsequent epidemiological interpretation.
@@
Since $(A^{k})_{ij}$ counts the number of walks of length $k$ between $i$ and $j$, communicability sums all possible routes, weighting longer paths by decaying factors $\beta^{k}/k!$. Thus, $C_{ij}(\beta)$ captures structural accessibility across multiple scales [7, 15, 29], from local neighborhoods to global connections.
@@ 4.3 Classical Spatial Models (SAR and SAC)
Traditional spatial models—the Spatial Autoregressive Model (SAR) and the Spatial Autoregressive Combined (SAC)—were fitted as an intermediate step to assess residual spatial dependence prior to hierarchical Bayesian modeling [29, 18]. Both models used the Queen contiguity matrix, despite the structural fragmentation of the mesh (33 disconnected components), a typical feature of cities with heterogeneous morphology [18].
Despite this topological limitation, the models produced stable estimates consistent with the remaining results. The SAR model exhibited Pseudo-$R^{2}=0.424$, while the SAC model yielded a similar value ($0.421$). These values indicate that a substantial portion of spatial variability is captured by the covariates, although enough residual autocorrelation remains to justify a hybrid spatial–random-effects approach (e.g., BYM2 or SPDE–INLA) [19].
@@
The spatial lag term $Wy$, however, was not statistically significant (SAR: $p=0.20$; SAC: $p=0.40$), suggesting that spatial dependence does not operate primarily through autocorrelation between adjacent areas, but rather through structural gradients in the urban environment—supporting the use of hierarchical models such as BYM2 [19, 32], which represent spatial smoothing more flexibly.
@@ 4.4 INLA/BYM2 Model: Latent Spatial Structure and Robustness
As the final stage of the empirical analysis, we estimated a hierarchical spatial model of the Besag–York–Mollié family in its BYM2 parametrization, using the Integrated Nested Laplace Approximation (INLA) [31, 19]. This model is widely employed in contemporary spatial epidemiology because it decomposes variability into a structured component (neighborhood-based) and an unstructured component, while providing accurate inference for the latent process via hyperparameters.
@@
Table 2: Fixed effects from the INLA/BYM2 model. Coefficients represent posterior means and 95% credible intervals.
Variable Mean SD 2.5% Median 97.5%
Structural and climatic covariates
Year (numeric) -11.75 0.90 -13.52 -11.75 -9.98
Curvature -46.41 3.99 -54.24 -46.41 -38.58
Mean annual precipitation -0.02 0.01 -0.04 -0.02 -0.004
Precipitation (lag 3) 0.37 0.07 0.24 0.37 0.50
Rainy season indicator 21.58 3.32 15.08 21.58 28.09
Socioeconomic covariates
Slum proportion (%) -25.84 12.93 -51.19 -25.85 -0.48
Population density 0.0056 0.0004 0.0048 0.0056 0.0065
Richest income share (%) -28.43 8.34 -44.79 -28.43 -12.06
Total population -0.00044 0.00015 -0.00074 -0.00044 -0.00014
@@
Table 3: BYM2 hyperparameters estimated using INLA.
Hyperparameter Mean SD 2.5% Median 97.5%
Precision (Gaussian) $1.39\cdot 10^{-4}$ $5.77\cdot 10^{-6}$ $1.28\cdot 10^{-4}$ $1.39\cdot 10^{-4}$ $1.51\cdot 10^{-4}$
@@
$\phi$ (structured share) $5.41\cdot 10^{-10}$ $1.16\cdot 10^{-10}$ $3.02\cdot 10^{-10}$ $5.48\cdot 10^{-10}$ $7.26\cdot 10^{-10}$
@@
These results indicate that the geometry of the urban network—not mere geographic proximity—structures dengue diffusion in Recife [26, 27].
@@ 6 Conclusion
7 Data and Code Availability
The geocoded dengue dataset used in this study is publicly available on Zenodo [22].
All scripts used for data cleaning, graph construction, communicability curvature computation, spatial modeling (Negative Binomial, SAR/SAC, and INLA/BYM2), and figure generation are available in an open GitHub repository [23].